# Calculating Bayesian evidence for inflationary models using CONNECT

### v1

@@ Abstract
Bayesian evidence is a standard method used for comparing the ability of different models to fit available data and is used extensively in cosmology. However, since the evidence calculation involves performing an integral of the likelihood function over the entire space of model parameters this can be prohibitively expensive in terms of both CPU and time consumption. For example, in the simplest $\Lambda$CDM model and using CMB data from the Planck satellite, the dimensionality of the model space is over 30 (typically 6 cosmological parameters and 28 nuisance parameters). Even the simplest possible model requires $\mathcal{O}(10^{6})$ calls to an Einstein–Boltzmann solver such as class or camb and takes several days.
@@ 1 Introduction
A way to mitigate this could be to use a cosmological emulator instead of the Einstein–Boltzmann solver code. Recent years has seen a surge in popularity of such emulators and they have been applied in many different ways. The most common kinds of emulators are either based on Artificial Neural Networks [17, 18, 19, 20] or Gaussian Processes [21, 22], both with their respective advantages and drawbacks. The applications range from standard Bayesian marginalisation to frequentist profile likelihoods [23], and Refs. [24, 25, 26] furthermore employed emulators to approximate Bayesian evidence using posterior samples and a modification of the harmonic mean estimator [27]. While approximations of the Bayesian evidence are useful to roughly compare cosmological models with very different evidence, models that only differ slightly need better estimates (e.g. from nested sampling) in order to perform a meaningful comparison.
@@
Finally, since we are using inflationary model selection as our test case, we must of course credit the pioneering work in Refs. [29, 30]. (See also Ref. [31] for a very recent update.) In these papers, the authors computed an effective likelihood by integrating out all non-inflationary parameters. A neural network was then trained to emulate this effective likelihood, allowing the authors to perform an exhaustive Bayesian model-comparison of most slow-roll inflationary models in the $\Lambda\text{CDM}$-model.
@@ 2 The connect framework and Bayesian evidence
The connect framework for emulation of cosmological observables has been tested extensively for cosmological parameter inference, using both Bayesian marginalisation and frequentist profile likelihoods [18, 23]. connect trains a neural network based on training data sampled iteratively to best represent the likelihood function. This ensures that the neural network is most precise where the likelihood is large, which makes it ideal for parameter inference. The training data is gathered using the fast Planck Lite likelihood [32]. The reason is that training requires a very large number of likelihood evaluations, which in the case of the full Planck likelihood would be prohibitively expensive. Because Planck Lite is somewhat less constraining than the full Planck likelihood, this gives us a set of training data that is more widely spread, and this along with a high sampling temperature yields a set of training data that can accurately represent several combinations of cosmological data sets (as long as either the full Planck likelihood, Planck Lite or similar CMB data is included) without the need to retrain the network.
@@
The Bayesian evidence has been calculated with the code PolyChord [12, 34], which uses a version of nested sampling [10] to calculate the evidence. The code is run from within the MCMC sampler MontePython [3, 35] with either class or connect as the cosmological theory code.
@@ 3 Validation of connect for evidence computation
$100\times\omega_{\mathrm{b}}$ $2.2$ $2.5$
@@
Since our main goal is to demonstrate the feasibility of using the connect framework for Bayesian evidence calculations, we will use the same data combination as in Ref. [28]. This will facilitate a more straightforward comparison between our results and those in Ref. [28]. Our data sets therefore in all cases consist of the full Planck 2018 TT,TE,EE+lowE data [32], the Planck 2018 lensing data [36], as well as the BICEP Keck 2015 data [37].
@@
Case Bayesian evidence ($\log\mathcal{Z}$) Number of likelihood calls
PolyChord with class $-1860.4\pm 0.54$ $1,419,152/115,988,382$
PolyChord with connect $-1861.1\pm 0.55$ $1,569,491$
Table 2: The Bayesian evidence calculated with PolyChord using both connect and class. The last column gives the total number of likelihood calls, in the case of the class-based run divided between cosmological and nuisance parameters.
@@
Figure 1: The posteriors for the physical and inflationary parameters from the calculation of the Bayesian evidence. The contours correspond to 68.3%, 95.5%, and 99.7% credible intervals. These posteriors are computed as a byproduct of the evidence calculation and are not expected to be as accurate as posteriors obtained through standard MCMC methods.
@@ 5 Numerical results
aspic model $\ln\mathcal{B}$ $\ln\mathcal{B}$ in Ref. [28]
Large Field Inflation (LFI2) $-8.8\pm 0.9$ $-11.5$
Large Field Inflation (LFI4) $-51.2\pm 0.9$ $-56.0$
Natural Inflation (NI) $-4.6\pm 0.9$ $-6.6$
Loop Inflation (LI) $-4.7\pm 0.9$ $-6.8$
Colemann-Weinberg Inflation (CWI) $-19.7\pm 1.0$ Not in the reference
Table 4: The calculated Bayesian evidence of the inflationary models with respect to the calculated Bayesian evidence for Higgs inflation. The uncertainties on the values from Ref. [28] is quoted as 0.3 in the article using 512 live points (note that estimated statistical uncertainties are typically significantly smaller for the same number of live points when using class because of the very large number of likelihood evaluations in the nuisance parameter space).
@@
Comparing our results with Ref. [28], it can be seen from Table 4 that they are in qualitative agreement regarding which models that are strongly disfavoured compared to Higgs Inflation. We note that the values for the Bayes factor found in this article do not match those from Ref. [28] completely, but this does not change the conclusion that all models are strongly disfavoured compared to Higgs Inflation. Furthermore, based on the tests performed in Section 3 and Appendix A, we are confident that connect gives results significantly closer to class than the differences observed here.
@@ 7 Discussion and conclusions
Based on the tests carried out and presented here we are therefore confident that connect can be used for calculations of Bayesian evidence in cosmology, vastly reducing the often prohibitive runtimes of such calculations.
Reproducibility. We have used the publicly available connect framework available at https://github.com/AarhusCosmology/connect_public to create training data and train neural networks. To calculate the Bayesian evidence, we have used the publicly available program PolyChord available at https://github.com/PolyChord/PolyChordLite as well as the program MontePython publicly available at https://github.com/brinckmann/montepython_public. Lastly, we have used the program aspic to calculate the inflationary models and their fundamental parameters. This has been done with the publicly available Python wrapper Pyaspic available at https://github.com/AarhusCosmology/PyAspic. aspic is publicly available at http://cp3.irmp.ucl.ac.be/~ringeval/upload/patches/aspic/.
@@ Acknowledgements
In this appendix we validate the use of PolyChord with connect and discuss convergence in terms of the number of live points. As we discussed in Section 3, the default setting for PolyChord in MontePython leads to severe undersampling of the nuisance parameter space when using connect  [^4] There are other situations where the default behaviour can be sub-optimal, see e.g. the issue at https://github.com/brinckmann/montepython_public/issues/374.. This leads to a bias in the evidence towards larger values as shown in Figure 2 for the $\text{LFI}_{4}$ inflationary model. We have performed 10 PolyChord runs for each setting of 300, 1200, 2400, and 4800 live points and from these it is quite evident that the evidence is strongly biased with these settings unless a very large number of live points is used.
Figure 2: The evidence for the inflation model LFI4 calculated ten times with respectively 300, 1200, 2400, and 4800 live points with connect and the default settings of PolyChord and MontePython. The scatter of the evidence is clearly much larger than the statistical variance reported by PolyChord, and the result is biased towards larger values of the evidence. The result seems to be well converged when using 2400 – 4800 live points.

### v2

@@ Abstract
Bayesian evidence is a standard tool used for comparing the ability of different models to fit available data and is used extensively in cosmology. However, since the evidence calculation involves performing an integral of the likelihood function over the entire space of model parameters this can be prohibitively expensive in terms of both CPU and time consumption. For example, in the simplest $\Lambda$CDM model and using CMB data from the Planck satellite, the dimensionality of the model space is over 30 (typically 6 cosmological parameters and 28 nuisance parameters). Even the simplest possible model requires $\mathcal{O}(10^{6})$ calls to an Einstein–Boltzmann solver such as class or camb and takes several days.
@@ 1 Introduction
A way to mitigate this could be to use a cosmological emulator instead of the Einstein–Boltzmann solver code. Recent years has seen a surge in popularity of such emulators and they have been applied in many different ways. The most common kinds of emulators are either based on Artificial Neural Networks [17, 18, 19, 20] or Gaussian Processes [21, 22], both with their respective advantages and drawbacks. The applications range from standard Bayesian marginalisation to frequentist profile likelihoods [23], and Refs. [24, 25, 26] furthermore employed emulators to approximate Bayesian evidence using posterior samples and a modification of the harmonic mean estimator [27]. While approximations of the Bayesian evidence are useful to roughly compare cosmological models with very different evidence, models that only differ slightly need better estimates (e.g. from nested sampling) in order to perform a meaningful comparison. Ref. [17] demonstrated that evidence computations could indeed be accurately computed using an emulator, albeit for the vanilla $\Lambda$CDM-model and using large-scale structure likelihoods and/or Planck Lite.
@@
Finally, since we are using inflationary model selection as our test case, we must of course credit the pioneering work in Refs. [29, 30]. (See also Ref. [31] for a very recent update.) In these papers, the authors computed an effective likelihood by integrating out all non-inflationary parameters. A neural network was then trained to emulate this effective likelihood, allowing the authors to perform an exhaustive Bayesian model-comparison of most slow-roll inflationary models in the $\Lambda\text{CDM}$-model. Furthermore, we must also mention the early work done on inflationary model selection in Ref. [32].
@@ 2 The connect framework and Bayesian evidence
The connect framework for emulation of cosmological observables has been tested extensively for cosmological parameter inference, using both Bayesian marginalisation and frequentist profile likelihoods [18, 23]. A main advantage of connect is that it contains both an emulator of cosmological observables as well as the framework needed to build and train the network. This means that emulators of new and non-standard cosmological models can easily be built and used to run cosmological parameter analyses within a single environment. connect trains a neural network based on training data sampled iteratively to best represent the likelihood function. This ensures that the neural network is most precise where the likelihood is large, which makes it ideal for parameter inference. The training data is gathered using the fast Planck Lite likelihood [33]. The reason is that training requires a very large number of likelihood evaluations, which in the case of the full Planck likelihood would be prohibitively expensive. Because Planck Lite is somewhat less constraining than the full Planck likelihood, this gives us a set of training data that is more widely spread and this (along with a high sampling temperature, which guarantees adequate coverage of the training data set), yields a set of training data that can accurately represent several combinations of cosmological data sets (as long as either the full Planck likelihood, Planck Lite or similar CMB data is included) without the need to retrain the network. Furthermore, connect is both the emulator as well as the procedure that trains the emulator. This makes it easy to quickly train a new physics model.
@@
The Bayesian evidence has been calculated with the code PolyChord [12, 35], which uses a version of nested sampling [10] to calculate the evidence. The code is run from within the MCMC sampler MontePython [3, 36] with either class or connect as the cosmological theory code. We finally note that it is only necessary to train one model with connect, because all the inflationary parameters for the different inflationary models can be mapped to the same physical parameters.
@@ 3 Validation of connect for evidence computation
$100\times\omega_{\mathrm{b}}$ $1.9$ $2.5$
@@
Since our main goal is to demonstrate the feasibility of using the connect framework for Bayesian evidence calculations, we will use the same data combination as in Ref. [28]. However, it should be stressed that evidence calculations are notoriously hard to compare because exact numbers are extremely sensitive to hyperparameter choices such as e.g. prior volume. Therefore, it is not to be expected that a direct, quantitative comparison can be made between our results and those of [28]. Our data sets therefore in all cases consist of the full Planck 2018 TT,TE,EE+lowE data [33], the Planck 2018 lensing data [37], as well as the BICEP Keck 2015 data [38].
@@
Case $\log\mathcal{Z}$ Likelihood calls CPU hours used
PolyChord with class $-1860.4\pm 0.54$ $1,419,152/115,988,382$ 30,000
PolyChord with connect $-1861.1\pm 0.55$ $1,569,491$ 125
Table 2: The Bayesian evidence calculated with PolyChord using both connect and class. 300 live points were used, and the oversampling feature was turned off for connect. Note that the log-evidence and the error estimates are comparable despite the class-run using $\sim 100$ times more likelihood calls. The last column shows the CPU core-hours used in each PolyChord run.
@@
Figure 1: The posteriors for the physical and inflationary parameters with the bounds given in Table 1. The contours correspond to 68.3%, 95.5%, and 99.7% credible intervals.
@@ 5 Numerical results
aspic model $\ln\mathcal{B}$
Large Field Inflation (LFI2) $-8.8\pm 0.4$
Large Field Inflation (LFI4) $-51.2\pm 0.4$
Natural Inflation (NI) $-4.6\pm 0.4$
Loop Inflation (LI) $-4.7\pm 0.4$
Colemann-Weinberg Inflation (CWI) $-19.7\pm 0.6$
Table 4: The calculated Bayesian evidence of the inflationary models with respect to the calculated Bayesian evidence for Higgs inflation. In [28], corresponding values for LFI2, LFI4, NI, and LI are -11.5, -56.0, -6.6, and -6.8 respectively. Coleman-Weinberg inflation was not tested in that work. The uncertainties on the values from Ref. [28] is quoted as 0.3 in the article using 512 live points (note that estimated statistical uncertainties are typically significantly smaller for the same number of live points when using class because of the very large number of likelihood evaluations in the nuisance parameter space).
@@
Comparing our results with Ref. [28], we find that they are in qualitative agreement regarding which models that are strongly disfavoured compared to Higgs Inflation. We note that the values for the Bayes factor found here are systematically more negative than the corresponding values in [28], and in most cases deviate more than the estimated statistical uncertainty. We stress that this is not due to problems related to the connect emulation, but is most likely related to small differences in the prior volume and/or parameterisation.
@@ 7 Discussion and conclusions
Based on the tests carried out and presented here we are therefore confident that connect can be used for calculations of Bayesian evidence in cosmology, vastly reducing the often prohibitive runtimes of such calculations. This will make it possible to start using Bayesian evidence as a tool in theoretical cosmology. Right now, theoretical cosmology is mostly done using Bayesian and frequentist parameter estimation, but it will now be possible to also use Bayesian evidence and thereby compare how good one cosmological model is compared to another cosmological model.
Reproducibility. We have used the publicly available connect framework available at https://github.com/AarhusCosmology/connect_public/tree/2406.03968 to create training data and train neural networks. To calculate the Bayesian evidence, we have used the publicly available program PolyChord available at https://github.com/PolyChord/PolyChordLite as well as the program MontePython publicly available at https://github.com/AarhusCosmology/montepython_public/tree/2406.03968. Lastly, we have used the program aspic to calculate the inflationary models and their fundamental parameters. This has been done with the publicly available Python wrapper Pyaspic available at https://github.com/AarhusCosmology/PyAspic. aspic is publicly available at http://cp3.irmp.ucl.ac.be/~ringeval/upload/patches/aspic/.
@@ Acknowledgements
In this appendix we validate the use of PolyChord with connect and discuss convergence in terms of the number of live points. As we discussed in Section 3, the default setting for PolyChord in MontePython leads to severe undersampling of the nuisance parameter space when using connect  [^4] There are other situations where the default behaviour can be sub-optimal, see e.g. the issue at https://github.com/brinckmann/montepython_public/issues/374.. This undersampling leads to a bias in the ensemble mean log-evidence, unless a very large number of live points is being used, as shown in Figure 2.